T be a linear transformation on V and A and B be two ordered bases for V and A is the matrix of T relative to A and B is the matrix of T relative to B then prove that A and B are similar.


1
Expert's answer
2021-07-12T13:10:04-0400

Let "\mathfrak{A}=(e_1, e_2,e_3)", "\mathfrak{B}=(f_1, f_2, f_3)" be two bases for the vector space V.

The fact that "A=(a_{ij})" is the matrix of linear transformation T relative to the basis "\mathfrak{A}" means that:

"Te_1=a_{11}e_1+a_{21}e_2+a_{31}e_3"

"Te_2=a_{12}e_1+a_{22}e_2+a_{32}e_3"

"Te_3=a_{13}e_1+a_{23}e_2+a_{33}e_3"

The fact that "B=(b_{ij})" is the matrix of linear transformation T relative to the basis "\mathfrak{B}" means that:

"Tf_1=b_{11}f_1+b_{21}f_2+b_{31}f_3"

"Tf_2=b_{12}f_1+b_{22}f_2+b_{32}f_3"

"Tf_3=b_{13}f_1+b_{23}f_2+b_{33}f_3"

Let C be the change-of-basis matrix:

"e_1=c_{11}f_1+c_{21}f_2+c_{31}f_3=\sum\limits_{i=1}^{3}c_{i1}f_i"

"e_2=c_{12}f_1+c_{22}f_2+c_{32}f_3=\sum\limits_{i=1}^{3}c_{i2}f_i"

"e_3=c_{13}f_1+c_{23}f_2+c_{33}f_3=\sum\limits_{i=1}^{3}c_{i3}f_i"

In general, "e_j=\sum\limits_{i=1}^{3}c_{ij}f_i".

Then

"Te_j=\sum\limits_{i=1}^{3}c_{ij}Tf_i=\sum\limits_{i=1}^{3}c_{ij}\sum\limits_{k=1}^{3}b_{ki}f_k=\sum\limits_{k=1}^{3}f_k\sum\limits_{i=1}^{3}c_{ij}b_{ki}"

From the other side

"Te_j=\sum\limits_{i=1}^{3}a_{ij}e_i=\sum\limits_{i=1}^{3}a_{ij}\sum\limits_{k=1}^{3}c_{ki}f_k=\sum\limits_{k=1}^{3}f_k\sum\limits_{i=1}^{3}a_{ij}c_{ki}"

Comparing these two formulas, we obtain

"\sum\limits_{i=1}^{3}c_{ij}b_{ki}=\sum\limits_{i=1}^{3}a_{ij}c_{ki}"

In matrices: BC=CA or "B=CAC^{-1}". This means that the matrices A and B are similar.